OpenStudy (anonymous):

Janet smooth pays her workers 175,495, and710 respectively per week. Calculate what Janet will pay t the end of the first quarter for a state unemployment and federal unemployment. Assume a state rate of 5.6% and federal .8% base is 7000

OpenStudy (anonymous):

I think you might need to use the A = Pe^rt equation for this

OpenStudy (tkhunny):

No, this is not interest accumulation. Quarterly Earnings 175 * 13 = 2275 495 * 13 = 6435 710 * 13 = 9230 State Rate: 2275 * 0.056 = 127.40 6435 * 0.056 = 360.36 9230 * 0.056 = 516.88 Federal Base Rate Rate: 2275 < 7000, so 2275 * 0.008 = 18.20 6435 < 7000, so 6435 * 0.008 = 51.48 9230 > 7000, so 7000 * 0.008 = 56.00 Just focus and wade through it. Spreadsheets can be wonderful things.

OpenStudy (anonymous):

lol yes they can, glad you found what you were looking for